# Heterogeneous Multi-Agent Proximal Policy Optimization for Power Distribution System Restoration

### v4

@@ Abstract
Restoring power distribution systems (PDS) after large-scale outages requires sequential switching operations that reconfigure feeder topology and coordinate distributed energy resources (DERs) under nonlinear constraints such as power balance, voltage limits, and thermal ratings. These challenges make conventional optimization and value-based RL approaches computationally inefficient and difficult to scale. This paper applies a Heterogeneous-Agent Reinforcement Learning (HARL) framework, instantiated through Heterogeneous-Agent Proximal Policy Optimization (HAPPO), to enable coordinated restoration across interconnected microgrids. Each agent controls a distinct microgrid with different loads, DER capacities, and switch counts, introducing practical structural heterogeneity. Decentralized actor policies are trained with a centralized critic to compute advantage values for stable on-policy updates. A physics-informed OpenDSS environment provides full power-flow feedback and enforces operational limits via differentiable penalty signals rather than invalid action masking. The total DER generation is capped at 2400 kW, and each microgrid must satisfy local supply–demand feasibility. Experiments on the IEEE 123-bus and IEEE 8500-node systems show that HAPPO achieves faster convergence, higher restored power, and smoother multi-seed training than DQN, PPO, MAES, MAGDPG, MADQN, Mean-Field RL, and QMIX. Results demonstrate that incorporating microgrid-level heterogeneity within the HARL framework yields a scalable, stable, and constraint-aware solution for complex PDS restoration.
@@ I Introduction
Modern power distribution systems face increasing exposure to environmental hazards, component aging, and cybersecurity threats, resulting in frequent service interruptions [1]. Distributed energy resources (DERs) such as photovoltaics, microturbines, and storage can accelerate local recovery [2], yet their limited capacity and the voltage, thermal, and topological constraints of distribution feeders complicate safe and efficient restoration [3]. Recent advancements in machine learning for critical infrastructure highlight the growing role of data-driven recovery, including a transformer–reinforcement learning framework for large-scale repair sequencing that enables faster and more resilient system restoration [4]. Machine learning has also advanced modeling in other safety-critical physical domains, including high-fidelity ground-motion prediction using ensemble techniques [5].
@@
Conventional restoration is typically formulated as a mixed-integer nonlinear optimization problem [6, 7, 8], which may provide optimal switching plans but requires accurate models and becomes computationally prohibitive for large or dynamically changing feeders [9]. Moreover, restoration is inherently sequential and stochastic, with each switching action altering future feasibility [10]. Reinforcement learning (RL) provides an alternative by learning control policies directly through interaction [11], supported by broader advances in deep and graph-based learning for power systems [12]. Early RL efforts used single-agent controllers based on DQN [13] or actor–critic methods [14], but these approaches suffered from biased value estimates [15], inefficient exploration, and limited scalability in feeders with many controllable devices [16].
@@
Multi-agent reinforcement learning (MARL) addresses scalability by assigning local controllers to network subregions. Methods such as MAES [17], MAGDPG [18], Mean-Field RL [19], and QMIX [20] improve coordination but still face key limitations: (i) reliance on value-based updates or fixed exploration rules, which amplify approximation errors as agents learn simultaneously; (ii) feasibility enforcement via action masking or early termination, which disrupts learning and leads to unsafe or infeasible actions; and (iii) overfitting to common operating patterns, reducing performance under unseen conditions.
@@
To overcome these challenges, this paper applies the Heterogeneous-Agent Proximal Policy Optimization (HAPPO) algorithm within a heterogeneous-agent reinforcement learning (HARL) architecture [22] for power distribution system restoration. HAPPO extends PPO to multi-agent settings through sequential per-agent updates guided by a centralized critic, enabling stable and sample-efficient learning among microgrids with differing load levels, generation capacities, and switch counts. These structural differences make agents non-exchangeable and unsuitable for parameter-sharing methods such as MAPPO.
@@
To enable physics-informed, constraint-aware learning, we develop an OpenDSS-based simulation environment [23] that performs full three-phase power-flow analysis after each switching action. Rather than masking infeasible actions [24] or terminating episodes [15], the environment applies a differentiable penalty term proportional to voltage, current, and generation-limit violations, while enforcing a 2400,kW system-level generation cap and local microgrid balance constraints. This soft-constraint design provides dense feedback and allows agents to learn corrective behavior under infeasible conditions.
@@
Experiments on the IEEE 123-bus [25] and IEEE 8500-node [26] feeders show that the proposed HAPPO framework achieves faster convergence, higher restored power, and improved stability compared with established MARL baselines including DQN [13], PPO [27], MAES [17], MAGDPG [18], MADQN [24], Mean-Field RL [19], and QMIX [20]. The contributions of this work are: (1) the first application of HAPPO to feeder-level restoration with structurally heterogeneous microgrid agents; (2) a stable and scalable learning framework combining decentralized actors with a centralized advantage critic; (3) a physics-informed OpenDSS environment with continuous differentiable constraint handling; and (4) extensive evaluation on IEEE test feeders demonstrating superior restoration performance and reproducibility.
@@ II-A Network Model and Decision Variables
Let the distribution network be represented by a graph $G=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}$ and $\mathcal{E}$ denote buses and branches, respectively. Loads lie on $\mathcal{D}\subseteq\mathcal{N}$ and distributed energy resources (DERs) on $\mathcal{G}\subseteq\mathcal{N}$. A subset $\mathcal{S}\subseteq\mathcal{E}$ contains controllable sectionalizing or tie switches, each with binary state $b_{q}\in\{0,1\}$ indicating a closed (energized) or open (de-energized) line. After an outage, restoration proceeds through discrete switching actions that reconfigure the topology and determine the energized buses and restored loads.
@@ II-B Scenario Randomization and Priorities
Each simulation episode introduces variability by randomizing the faulted or unavailable branches, the placement of critical loads, and any optional changes in DER penetration levels.
Each load $k\in\mathcal{D}$ is assigned a priority weight $c_{k}\in\{1,\dots,10\}$, indicating its relative importance in the restoration process. These priority assignments differ between feeders and determine the weighted-load objective defined below.
II-C Restoration Objective
@@
II-D Operational Constraints
At each switching configuration, the distribution system must satisfy fundamental electrical and operational limits.
II-D1 Power Balance
The total restored load and network losses must not exceed the available DER generation:
$\sum_{k\in\mathcal{D}}P_{k}+\sum_{\ell\in\mathcal{E}}P^{\text{loss}}_{\ell}\leq\sum_{g\in\mathcal{G}}P_{g}.$ (C1)
II-D2 Voltage Limits
@@
$V_{\min}\leq V_{i}\leq V_{\max},\quad\forall i\in\mathcal{N}.$ (C2)
II-D3 DER Operating Limits
@@
$P_{g}^{\min}\leq P_{g}\leq P_{g}^{\max},\quad Q_{g}^{\min}\leq Q_{g}\leq Q_{g}^{\max},\quad\forall g\in\mathcal{G}.$ (C3)
II-D4 Thermal Line Limits
@@
$P_{ij}^{2}+Q_{ij}^{2}\leq(S^{\max}_{ij})^{2},\quad\forall(i,j)\in\mathcal{E}.$ (C4)
II-D5 Global Generation Capacity
@@
$\sum_{g\in\mathcal{G}}P_{g}\leq 2400~\text{kW}.$ (C5)
II-D6 Local Microgrid Balance
For each microgrid $m$, local load consumption cannot exceed its local DER generation:
$P^{\text{load}}_{m}\leq P^{\text{gen}}_{m},\quad\forall m.$ (C6)
These constraints define the feasible space of valid switching configurations during restoration.
II-E Structural Heterogeneity and Feeder Partitioning
To reduce the combinatorial complexity of the restoration problem, the feeder is partitioned into $N$ microgrids. Each microgrid includes its own subset of loads and DERs, operates a distinct set of controllable switches $\mathcal{Q}_{i}\subseteq\mathcal{S}$, and exhibits unique load magnitudes and generation capacities, resulting in structurally heterogeneous operating regions. Although all switches share the same action type (open/close/no-op), the number of switches, load profiles, DER availability, and topological location vary significantly across microgrids. This introduces structural heterogeneity, making microgrids non-exchangeable and motivating the heterogeneous-agent formulation adopted in later sections.
II-F Reward Function (for RL-Based Optimization)
When solving (1) via reinforcement learning, a dense reward function is constructed to encourage restoration progress while penalizing constraint violations $r_{t}=\alpha\,\Delta P^{\text{rest}}_{t}-\beta\frac{P^{\text{loss}}_{t}}{P_{\text{gen}}}-\lambda\,\xi_{t},$ where $\Delta P^{\text{rest}}_{t}$ represents the incremental increase in restored power, and $P_{\text{loss}}$ is normalized by the fixed generation cap $P_{\text{gen}}=2400~\text{kW}$. The penalty term $\xi_{t}$ collects the magnitudes of violations of constraints (C1)–(C6), whereas the parameters $\alpha$, $\beta$, and $\lambda$ serve as nonnegative weights that tune the contributions of restoration progress, loss minimization, and constraint enforcement, respectively.
This reward formulation preserves the mathematical objective while enabling learning-based optimization in Section III.
@@ III-A HARL Architecture Overview
The proposed approach is built within a heterogeneous-agent reinforcement learning (HARL) architecture [22], where each microgrid agent maintains an independent actor policy and a shared centralized critic estimates global value functions. This design reflects the physical hierarchy of distribution systems: local agents observe and control only their own microgrids, while the critic leverages system-wide information to provide coordinated learning signals.
Each agent $i$ follows a policy $a_{i,t}=\pi_{\theta_{i}}(o_{i,t}),$ issuing a discrete switching action (open, close, or no-op). The centralized critic $V_{\phi}(s_{t})$ processes the global system state to compute value and advantage estimates that guide all agents.
@@ III-B Agent–Environment Interaction
At each simulation step $t$, each agent receives a local observation $o_{i,t}$ (e.g., voltages, currents, DER outputs, and switch states), selects an action $a_{i,t}=\pi_{\theta_{i}}(o_{i,t})$, and the joint action $a_{t}=(a_{1,t},\dots,a_{N,t})$ is executed in the OpenDSS simulator. OpenDSS computes the resulting feeder state, including next global state $s_{t+1}$, updated local observations $\{o_{i,t+1}\}$, restored power and system losses, voltage and current magnitudes, and the shared global reward $r_{t}$.
Each transition contributes to an on-policy trajectory buffer $(o^{i}_{t},a^{i}_{t},r_{t},o^{i}_{t+1},\psi_{t}),$ where $\psi_{t}$ contains shared global signals such as $s_{t}$ or the joint action $a_{t}$. These trajectories support centralized advantage computation and sequential policy optimization.
@@ III-C Advantage Computation via $\lambda$-GAE
$\delta_{t}$ $=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t}),$ (2)
$\hat{A}^{i}_{t}$ $=\sum_{\ell=0}^{T-t-1}(\gamma\lambda)^{\ell}\,\delta_{t+\ell},$ (3)
where $\gamma$ is the discount factor and $\lambda\in[0,1]$ controls the bias–variance trade-off. The centralized critic provides low-variance, system-consistent value estimates, enabling each agent to compute advantages aligned with the global restoration objective.
@@ III-D Sequential Policy Updates (HAPPO)
$L_{i}(\theta_{i})=\mathbb{E}\!\left[\min\!\left(r_{i,t}\hat{A}^{i}_{t},\,\text{clip}(r_{i,t},1-\epsilon,1+\epsilon)\hat{A}^{i}_{t}\right)\right]-\beta_{\mathrm{ent}}\,\mathbb{E}[H(\pi_{\theta_{i}})],$ (4)
where the probability ratio is $r_{i,t}=\dfrac{\pi_{\theta_{i}}(a_{i,t}\mid o_{i,t})}{\pi^{\text{old}}_{\theta_{i}}(a_{i,t}\mid o_{i,t})}.$
The clipping term enforces a trust-region constraint, preventing large, destabilizing policy shifts, while the entropy term $\beta_{\mathrm{ent}}$ encourages adequate exploration. Sequential updates reduce gradient interference and support monotonic policy improvement.
Critic update: The critic parameters $\phi$ are trained by regressing $V_{\phi}(s_{t})$ onto bootstrapped returns:
$L_{\text{critic}}=\mathbb{E}\!\left[(V_{\phi}(s_{t})-\hat{R}_{t})^{2}\right],\qquad\hat{R}_{t}=\hat{A}_{t}+V_{\phi}(s_{t}).$ (5)

### v5

@@ Abstract
Restoring power distribution systems (PDSs) after large-scale outages requires sequential switching actions that reconfigure feeder topology and coordinate distributed energy resources (DERs) under nonlinear constraints, including power balance, voltage limits, and thermal ratings. These challenges limit the scalability of conventional optimization and value-based reinforcement learning (RL) approaches. This paper applies a Heterogeneous-Agent Reinforcement Learning (HARL) framework via Heterogeneous-Agent Proximal Policy Optimization (HAPPO) to enable coordinated restoration across interconnected microgrids. Each agent controls a distinct microgrid with different loads, DER capacities, and switch counts. Decentralized actors are trained with a centralized critic for stable on-policy learning, while a physics-informed OpenDSS environment enforces electrical feasibility. Experiments on IEEE 123-bus and 8500-node feeders show HAPPO outperforms PPO, QMIX, Mean-Field RL, and other baselines in restored power, convergence stability, and multi-seed reproducibility. Under a 2400 kW generation cap, the framework restores over 95% of available load on both systems with low-latency execution, supporting practical real-time PDS restoration.
@@ I Introduction
Modern power distribution systems face increasing exposure to environmental hazards, component aging, and cybersecurity threats, resulting in frequent service interruptions [1]. Distributed energy resources (DERs) such as photovoltaics, microturbines, and storage can accelerate local recovery [2], yet their limited capacity and the voltage, thermal, and topological constraints of distribution feeders complicate safe and efficient restoration [3]. Recent advancements in machine learning for critical infrastructure highlight the growing role of data-driven recovery, including transformer–reinforcement learning frameworks for large-scale repair sequencing [4] and ensemble techniques for high-fidelity ground-motion prediction [5].
@@
Conventional restoration is typically formulated as a mixed-integer nonlinear optimization problem [6, 7, 8], which may provide optimal switching plans but becomes computationally prohibitive for large or dynamically changing feeders [9]. Moreover, restoration is inherently sequential and stochastic, with each switching action altering future feasibility [10]. Reinforcement learning (RL) provides an alternative by learning control policies directly through interaction [11], supported by broader advances in deep and graph-based learning for power systems [12]. Early RL efforts used single-agent controllers based on DQN [13] or actor–critic methods [14], but suffered from biased value estimates [15], inefficient exploration, and limited scalability [16].
@@
Multi-agent reinforcement learning (MARL) addresses scalability by assigning local controllers to network subregions. Methods such as MAES [17], MAGDPG [18], Mean-Field RL [19], and QMIX [20] improve coordination but face key limitations: (i) reliance on value-based updates or fixed exploration rules, amplifying approximation errors; (ii) feasibility enforcement via action masking or early termination, disrupting learning; and (iii) overfitting to common operating patterns.
@@
To overcome these challenges, this paper applies Heterogeneous-Agent Proximal Policy Optimization (HAPPO) within a heterogeneous-agent reinforcement learning (HARL) architecture [22] for power distribution system restoration. HAPPO extends PPO to multi-agent settings through sequential per-agent updates guided by a centralized critic, enabling stable learning among microgrids with differing load levels, generation capacities, and switch counts. These structural differences make agents non-exchangeable and unsuitable for parameter-sharing methods such as MAPPO.
@@
We develop an OpenDSS-based simulation environment [23] performing full three-phase power-flow analysis after each switching action. Rather than masking infeasible actions [24] or terminating episodes [15], the environment applies differentiable penalty terms proportional to voltage, current, and generation-limit violations, while enforcing a system-level generation cap and local microgrid balance constraints. This soft-constraint design provides dense feedback and allows agents to learn corrective behavior.
@@
Experiments on IEEE 123-bus [25] and 8500-node [26] feeders show the proposed HAPPO framework achieves faster convergence, higher restored power, and improved stability compared with DQN [13], PPO [27], MAES [17], MAGDPG [18], MADQN [24], Mean-Field RL [19], and QMIX [20]. The contributions are: (1) first application of HAPPO to feeder-level restoration with structurally heterogeneous microgrid agents; (2) a stable and scalable learning framework combining decentralized actors with a centralized advantage critic.
@@ II-A Network Model and Decision Variables
Let the distribution network be represented by a graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}$ denotes the set of buses and $\mathcal{E}$ denotes the set of branches. Loads lie on $\mathcal{D}\subseteq\mathcal{N}$ and distributed energy resources (DERs) on $\mathcal{G}\subseteq\mathcal{N}$. A subset $\mathcal{S}\subseteq\mathcal{E}$ contains controllable sectionalizing or tie switches. Each switch $q\in\mathcal{S}$ has a binary state $b_{q}\in\{0,1\}$ indicating a closed (energized) or open (de-energized) line. After an outage, restoration proceeds through discrete switching actions that reconfigure the topology and determine the energized buses and restored loads.
@@ II-B Scenario Randomization and Priorities
Each simulation episode introduces variability by randomizing the faulted or unavailable branches, the placement of critical loads, and any optional changes in DER penetration levels. Each load $k\in\mathcal{D}$ is assigned a priority weight $c_{k}\in\{1,\dots,10\}$, indicating its relative importance in the restoration process. These priority assignments differ between feeders and determine the weighted-load objective defined below.
II-C Restoration Objective and Operational Constraints
@@
At each switching configuration, the distribution system must satisfy fundamental electrical and operational limits. These constraints define the feasible space of valid switching configurations during restoration.
II-C1 Power Balance
The total restored load and network losses must not exceed the available DER generation. Here, the inequality permits unused generation capacity during partial restoration:
$\sum_{k\in\mathcal{D}}P_{k}+\sum_{e\in\mathcal{E}}P^{\text{loss}}_{e}\leq\sum_{g\in\mathcal{G}}P_{g},$ (2)
where $P_{k}$ is the active power consumed by load $k$, $P^{\text{loss}}_{e}$ is the power loss on branch $e$, and $P_{g}$ is the active power generated by DER unit $g$.
II-C2 Voltage Limits
@@
$V_{\min}\leq V_{k}\leq V_{\max},\quad\forall k\in\mathcal{N},$ (3)
where $V_{k}$ is the voltage magnitude at bus $k$.
II-C3 DER Operating Limits
@@
$P_{g}^{\min}\leq P_{g}\leq P_{g}^{\max},\quad Q_{g}^{\min}\leq Q_{g}\leq Q_{g}^{\max},\quad\forall g\in\mathcal{G},$
where $P_{g}$ and $Q_{g}$ are the active and reactive power outputs of DER unit $g$, respectively.
II-C4 Thermal Line Limits
@@
$P_{e}^{2}+Q_{e}^{2}\leq(S^{\max}_{e})^{2},\quad\forall e\in\mathcal{E},$ (5)
where $P_{e}$ and $Q_{e}$ are the active and reactive power flows on branch $e$, and $S^{\max}_{e}$ is the thermal capacity of branch $e$.
II-C5 Global Generation Capacity
@@
$\sum_{g\in\mathcal{G}}P_{g}\leq 2400~\text{kW}.$ (6)
II-C6 Local Microgrid Balance
For each microgrid $i$, local load consumption cannot exceed its local DER generation:
$P^{\text{load}}_{i}\leq P^{\text{gen}}_{i},\quad\forall i.$ (7)
II-D Structural Heterogeneity and Feeder Partitioning
To reduce the combinatorial complexity of the restoration problem, the feeder is partitioned into $N$ microgrids. Each microgrid $i$ includes its own subset of loads and DERs, operates a distinct set of controllable switches $\mathcal{Q}_{i}\subseteq\mathcal{S}$, and exhibits unique load magnitudes and generation capacities, resulting in structurally heterogeneous operating regions. While each agent controls switches with identical action types (open/close/no-op), microgrids differ substantially in the number of switches they control, their load profiles, DER availability, and topological structure. This structural heterogeneity makes microgrids non-exchangeable and motivates the heterogeneous-agent formulation adopted in later sections.
Microgrids are defined following standard feeder partitioning practices based on topology and DER locality, consistent with prior restoration benchmarks to ensure fair comparison. This choice reflects physical connectivity and operational coupling, and alternative reasonable partitions yielded similar convergence behavior, indicating low sensitivity to minor variations.
II-E Reward Function
When solving (1) via reinforcement learning, a dense reward function is constructed to encourage restoration progress while penalizing constraint violations. Let $\Delta P_{t}^{\text{rest}}=\sum_{k\in\mathcal{D}}c_{k}(P_{k,t}-P_{k,t-1})$ denote the weighted incremental increase in restored power at step $t$, where $P_{k,t}$ is the power restored to load $k$ at time $t$. Then $r_{t}=\alpha\,\Delta P^{\text{rest}}_{t}-\beta\frac{P^{\text{loss}}_{t}}{P_{\text{gen}}}-\lambda\,\xi_{t},$ where $\Delta P^{\text{rest}}_{t}$ represents the incremental increase in restored power, and $P^{\text{loss}}$ is normalized by the fixed generation cap $P_{\text{gen}}=2400~\text{kW}$. The penalty term $\xi_{t}$ collects the magnitudes of violations of constraints (2)–(7), whereas the parameters $\alpha$, $\beta$, and $\lambda$ serve as nonnegative weights that tune the contributions of restoration progress, loss minimization, and constraint enforcement, respectively. This reward formulation preserves the mathematical objective while enabling learning-based optimization in Section III.
@@ III-A HARL Architecture Overview
The proposed approach is built within a heterogeneous-agent reinforcement learning (HARL) architecture [22], where each microgrid agent maintains an independent actor policy and a shared centralized critic estimates global value functions. This design reflects the physical hierarchy of distribution systems: local agents observe and control only their own microgrids, while the critic leverages system-wide information to provide coordinated learning signals. Each agent $i$ follows a policy $a_{i,t}\sim\pi_{\theta_{i}}(o_{i,t}),$ issuing a discrete switching action (open, close, or no-op). The centralized critic $V_{\phi}(s_{t})$ processes the global system state to compute value and advantage estimates that guide all agents.
@@ III-B Agent–Environment Interaction
At each simulation step $t$, each agent $i$ receives a local observation $o_{i,t}$ (e.g., voltages, currents, DER outputs, and switch states), selects an action $a_{i,t}\sim\pi_{\theta_{i}}(o_{i,t})$, and the joint action $a_{t}=(a^{1}_{t},\ldots,a^{N}_{t})$ is executed in the OpenDSS simulator. OpenDSS computes the resulting feeder state, including next global state $s_{t+1}$, updated local observations $\{o_{i,t+1}\}$, restored power and system losses, voltage and current magnitudes, and the shared global reward $r_{t}$.
Each transition contributes to an on-policy trajectory buffer $(o^{i}_{t},a^{i}_{t},r_{t},o^{i}_{t+1},\psi_{t}),$ where $\psi_{t}$ contains auxiliary information such as the joint action $\mathbf{a}_{t}=(a_{1,t},\ldots,a_{N,t})$ and other global signals used for centralized training. These trajectories support centralized advantage computation and sequential policy optimization.
@@ III-C Advantage Computation via $\lambda$-GAE
$\delta_{t}^{i}$ $=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t}),$ (8)
$\hat{A}^{i}_{t}$ $=\sum_{\ell=0}^{T-t-1}(\gamma\lambda)^{\ell}\,\delta_{t+\ell}^{i},$ (9)
where $\gamma$ is the discount factor and $\lambda\in[0,1]$ controls the bias–variance trade-off, and $\ell$ is the time-step offset. Note that while the TD error uses the global reward $r_{t}$ and centralized value function $V_{\phi}(s_{t})$, advantages are computed separately for each agent $i$ to support sequential policy updates. The centralized critic provides low-variance, system-consistent value estimates, enabling each agent to compute advantages aligned with the global restoration objective.
@@ III-D Sequential Policy Updates (HAPPO)
$L_{i}(\theta_{i})=\mathbb{E}\!\left[\min\!\left(\rho_{i,t}\hat{A}^{i}_{t},\,\text{clip}(r_{i,t},1-\epsilon,1+\epsilon)\hat{A}^{i}_{t}\right)\right]-\beta_{\mathrm{ent}}\,\mathbb{E}[H(\pi_{\theta_{i}})],$
where the probability ratio is $\rho_{i,t}=\dfrac{\pi^{\text{new}}_{\theta_{i}}(a_{i,t}\mid o_{i,t})}{\pi^{\text{old}}_{\theta_{i}}(a_{i,t}\mid o_{i,t})}.$ The entropy regularization term $H(\pi_{\theta_{i}})$ encourages exploration, with $\beta_{\mathrm{ent}}$ controlling the strength of the entropy bonus.
The clipping term enforces a trust-region constraint, preventing large, destabilizing policy shifts, while the entropy term $\beta_{\mathrm{ent}}$ encourages adequate exploration. Sequential updates reduce gradient interference and support monotonic policy improvement. The critic parameters $\phi$ are trained by regressing $V_{\phi}(s_{t})$ onto bootstrapped returns:
$L_{\text{critic}}=\mathbb{E}\left[(V_{\phi}(s_{t})-\hat{R}_{t})^{2}\right],\quad\hat{R}_{t}=\sum_{\ell=0}^{T-t-1}\gamma^{\ell}r_{t+\ell}+\gamma^{T-t}V_{\phi}(s_{T}).$